# xbx-supp-arxiv-0

### v1

@@ Abstract
We introduce the XBX regression model, a continuous mixture of extended-support beta regressions for modeling bounded responses with or without boundary observations. The core building block of the new model is the extended-support beta distribution, which is a censored version of a four-parameter beta distribution with the same exceedance on the left and right of $(0,1)$. Hence, XBX regression is a direct extension of beta regression. We prove that both beta regression with dispersion effects and heteroscedastic normal regression with censoring at both 00 and $1$ — known as the heteroscedastic two-limit tobit model in the econometrics literature — are special cases of the extended-support beta regression model, depending on whether a single extra parameter is zero or infinity, respectively. To overcome identifiability issues that may arise in estimating the extra parameter due to the similarity of the beta and normal distribution for certain parameter settings, we assume that the additional parameter has an exponential distribution with an unknown mean. The associated marginal likelihood can be conveniently and accurately approximated using a Gauss-Laguerre quadrature rule, resulting in efficient estimation and inference procedures. The new model is used to analyze investment decisions in a behavioral economics experiment, where the occurrence and extent of loss aversion is of interest. In contrast to standard approaches, XBX regression can simultaneously capture the probability of rational behavior as well as the mean amount of loss aversion. Moreover, the effectiveness of the new model is illustrated through extensive numerical comparisons with alternative models.
@@ 1.1 Bounded responses and beta regression
From the available approaches, beta regression (Ferrari and Cribari-Neto, 2004; Smithson and Verkuilen, 2006) has emerged as the dominant likelihood-based approach for modeling bounded responses in terms of covariates. It has been successfully applied in diverse fields, including biology, ecology, economics and social sciences, finance, manufacturing, genetics, and engineering; see, for example, Allenbrand and Sherwood (2023), Leng et al. (2021), Namin et al. (2020), and Wang et al. (2015) for a few noteworthy applications, and Douma and Weedon (2019) and Geissinger et al. (2022) for recent articles making a case for its use in natural sciences and associated applied fields. The reason for its popularity is that beta regression respects the bounded range of the responses, and allows their mean and dispersion to be expressed in terms of covariates through separate regression structures. Such a construct offers high flexibility in the shape of the response distribution, capturing both the heteroscedasticity and the asymmetries observed in bounded responses.
@@
where $\mu\in(0,1)$, $\phi>0$, $B(p,q)$ $(q,p>0)$ is the beta function (see, Abramowitz and Stegun, 1964, §6.2.1), and $I(A)$ is the indicator function taking value $1$ if $A$ is true, and zero, otherwise. In this parameterization $\mathop{\rm E}(Y\mid\mu,\phi)=\mu$ and $\mathop{\rm var}(Y\mid\mu,\phi)=\mu(1-\mu)/(1+\phi)$ so that $\phi$ is a precision parameter. The beta regression model assumes that the observed responses $y_{1},\ldots,y_{n}$ are realizations of beta random variables $Y_{1},\ldots,Y_{n}$ that are independent conditionally on explanatory vectors $x_{1},\ldots,x_{n}$ and $z_{1},\ldots,z_{n}$, where $x_{i}$ and $z_{i}$ are $p$- and $q$-dimensional vectors, respectively, possibly sharing some or all components. As is apparent, we are assuming here that the observed responses take value between zero and one or have been transformed to do so as $y_{i}=(y_{i}^{\prime}-a)/(b-a)$, where $y_{i}^{\prime}$ are the original responses and $a$ and $b$ are the minimum and maximum value they can take. The mean and precision of $Y_{i}$ are then linked to the explanatory variables according to
@@
The parameter vectors $\beta$ and $\gamma$ are typically estimated by maximum likelihood, and standard inferential procedures, such as likelihood ratio, Wald, and score tests, apply on the grounds of central limit theorems for the derivatives of the log-likelihood under assumptions on model adequacy. Grün et al. (2012) provide more details and describe a range of modeling strategies and methods based on the beta regression model.
@@ 1.2 Boundary observations
One important limitation of the beta regression model as defined by (1) and (2) is that likelihood-based estimation and inference is not possible when at least one of the observed responses is either zero or one. Specifically, depending on the parameter values, the density (1) at those observations is either zero or infinity. As a result and regardless of the sample size, the log-likelihood of the beta regression model is either zero or infinite or undetermined, and maximum likelihood estimation and the associated inferential procedures cannot be used. This is the reason that popular software for beta regression modeling, such as the betareg R package (Cribari-Neto and Zeileis, 2010; Grün et al., 2012) prior to version 3.2-0, returns an error when there is at least one response value that is zero or one.
One popular and straightforward approach to deal with boundary observations is by Smithson and Verkuilen (2006), who propose to transform the response to $(y_{i}+u)/(1+2u)$, with $u=1/\{2(n-1)\}$, and then fit the beta regression model in (1) and (2) with standard software. The transformed responses, then lie in $(0.5/n,1-0.5/n)\subset(0,1)$, and, as a result, there are no issues when fitting the beta regression model. The proposal in Smithson and Verkuilen (2006) has the advantage of preserving the parsimony and the convenience in estimation and inference that is provided by beta regression model in (1) and (2), but the value of $u$ is rather ad hoc and, as Example 1 demonstrates, can have a marked effect on inference.
@@
Smithson and Verkuilen (2006, Example 3) models the contribution of nonverbal IQ and dyslexic versus nondyslexic status to the scores of 44 children on a test of reading accuracy. The scores were rescaled to lie in the unit interval, and 13 out of 44 children achieved a score of one. We transform the responses using $u=1/87$, and fit the beta regression model with $g_{1}(\mu_{i})=\log\{\mu_{i}/(1-\mu_{i})\}$, $g_{2}(\phi_{i})=\log(\phi_{i})$ in (2), where $\eta_{i}$ involves an intercept and the main and interaction effects of IQ and dyslexia status, and $\zeta_{i}$ involves only the main effects of IQ and dyslexia status. [^1] The numerical figures in Smithson and Verkuilen (2006, Table 5) can be reproduced by transforming only the boundary responses, and leaving the non-boundary ones unaltered. The same is true with Smithson and Merkle (2013, Table 6.2). We focus on the evidence against the omission of the interaction term.
Figure 1: Likelihood ratio and Wald statistics for the inclusion of an interaction term between dyslexia status and IQ in the setting of Example 1 versus $u$ ($\log_{10}$ scale). The vertical lines are the quantiles of the $\chi^{2}_{1}$ (likelihood ratio test) and normal distribution (two-sided Wald test) at levels $1.0\%$, $2.5\%$, $5.0\%$, $10.0\%$. The vertical dashed line is at $u=1/87$ as per the proposal in Smithson and Verkuilen (2006).
@@
$\mathop{\rm E}(Y)=P(Y^{*}\geq 1\mid\theta)+P(0<Y^{*}<1\mid\theta)\mathop{\rm E}(Y^{*}\mid 0<Y^{*}<1,\theta)$ (3)
@@
where $\prod_{A}a$ is $1$ if $A$ is the empty set. The standard approach towards defining a censored regression model is to put a regression structure on the components of $\theta$. The most widely-used censored regression model is the censored normal one, also known as the two-limit tobit. This model assumes that $Y^{*}_{1},\ldots,Y^{*}_{n}$ are independent conditionally on covariate vectors $x_{1},\ldots,x_{n}$, with $Y^{*}_{i}$ distributed according to a normal distribution with mean $x_{i}^{\top}\beta^{*}$ and variance $\sigma_{i}^{2}=\sigma^{2}$ (see, for example Rosett and Nelson, 1975; Maddala, 1983, Section 6.7, for definition and discussion). The heteroscedastic version of the two-limit tobit assumes that the variances are linked to explanatory variables as $g(\sigma_{i})=z_{i}^{\top}\gamma^{*}$, where $g(\cdot)$ is a monotone link function that maps $(0,\infty)$ to the real line (see, for example, Messner et al., 2016). Dixon and Sonka (1982) compare least squares and maximum likelihood estimation of the two-limit tobit in terms of estimates in an empirical application, and Hoff (2007) compares the two-limit tobit model to the semi-parametric approach in Papke and Wooldridge (1996) and the inflated beta distribution in Cook et al. (2008) for data envelopment analysis. Dixon and Sonka (1982), Hoff (2007), and Greene (2011, §19.3) touch on the interpretation of the parameters of the two-limit tobit. The disadvantage of the two-limit tobit is that it is less flexible than the beta distribution and cannot accomodate features that often occur in bounded-domain variables such as high skewness or kurtosis.
@@ 1.3 Our contribution
Section 2 presents the XB distribution and shows that the beta and the censored normal distribution are its limiting cases. Section 3 introduces the XBX regression model, and discusses parameter estimation and inference using approximate likelihood. Section 4 illustrates the benefits of XBX regression through the analysis of a loss aversion experiment from behavioral economics. The XBX model leads to qualitatively the same results for the effects about mean loss aversion as the original analysis in Glätzle-Rützler et al. (2015) but additionally allows for economically relevant insights about the probability of rational behavior. Section 5 presents a comprehensive simulation experiment comparing XBX regression to the heteroscedastic two-limit tobit model based on the continuous ranked probability score (CRPS). We conclude with remarks and directions for further research in Section 6.
@@ 2 Extended-support beta distribution
    If $\mu_{*}=\mathop{\rm E}(Y|\mu,\phi,u)\in\Re$ and $\sigma_{*}^{2}=\mathop{\rm var}(Y|\mu,\phi,u)\in(0,\infty)$ are fixed, then, as $u\to\infty$, $Y$ converges in distribution to $\mu_{*}+\sigma_{*}Z$, where $Z$ is a standard normal random variable.
@@
$h(t|\mu,\phi,u)=\frac{\sigma(\mu+\sigma t)^{\alpha-1}(1-\mu-\sigma t)^{\beta-1}}{B(\alpha,\beta)}I\left(\frac{-u-\mu_{*}}{\sigma_{*}}<t<\frac{1+u-\mu_{*}}{\sigma_{*}}\right)\,.$ (6)
@@
A direct corollary of Theorem 1 is that the XB distribution coincides with the beta distribution for $u=0$, and that it converges to the normal distribution censored in $[0,1]$ as $u\to\infty$. Hence, regression models based on the XB distribution allow to parameterize the transition from beta regression to the heteroscedastic two-limit tobit, at the expense of a single extra parameter $u$. The beta distribution and the normal distribution limits of the XB distribution can assume similar shapes in $(0,1)$. Hence, it is natural to consider shrinking $u$ to 00, or, equivalently, shrinking the XB distributions towards the beta distribution, to avoid any identifiability issues that may arise in the estimation of the parameters of the XB distribution, while preserving the flexibility offered by the beta distribution.
@@ 3.1 Exponential extended-support beta mixture
where $f_{\rm(XB)}(y\mid\mu,\phi,u)$ is defined in (7). By its definition, for fixed $\phi$ and $\nu$, the XBX density is reflection symmetric in that $f_{\rm(XBX)}(y\mid\mu,\phi,\nu)=f_{\rm(XBX)}(1-y\mid 1-\mu,\phi,\nu)$. Density (8) is not available in closed form, but it can be accurately and efficiently approximated using a Gauss-Laguerre quadrature rule (see, Abramowitz and Stegun, 1964, Equation 25.4.45) as
@@
Figure 3: Densities of the beta and XBX distribution for all combinations of $\mu\in\{0.5,0.7,0.9\}$, $\phi\in\{0.5,2,20\}$, and $\nu\in\{0.01,0.1,1\}$. The horizontal dotted line is at 1.
@@ 3.2 XBX regression
From expressions (10), the parameters $\mu$ and $\phi$ have direct interpretations in terms of location and precision of the distribution of $Y^{*}$, as they do in beta regression. In particular, the expectation of $Y^{*}$ is a linear function of $\mu$, with slope $(1+2\nu)$ and intercept $-\nu$. On the other hand, the variance of $Y^{*}$ decreases as $\phi$ increases, with the rate of decrease increasing quadratically with $\nu$. So, an XBX regression with mean and dispersion effects for the latent distribution can directly be defined by linking $\mu$ and $\phi$ to covariates as in (2). From Figure 3, it is directly apparent that the response distribution in XBX regression can have much more flexible shapes than the censored normal distribution underlying the heteroscedastic two-limit tobit; see, also Figure LABEL:fig-supp:xbx-vs-cn and Figure LABEL:fig-supp:xbx-vs-cn-hist in Section LABEL:sec-supp:xbx-vs-cn of the supplementary material document for a side-by-side comparison of the XBX distribution to the doubly-censored normal distribution.

### v2

# xbx-supp-arxiv-1

@@ Abstract
We introduce the XBX regression model, a continuous mixture of extended-support beta regressions for modelling bounded responses with boundary observations. The core building block of XBX regression is the extended-support beta distribution, a censored version of a four-parameter beta distribution with the same exceedance on the left and right of $(0,1)$. Hence, XBX regression is a direct extension of beta regression. We prove that beta regression and heteroscedastic normal regression with censoring at both 00 and $1$ — also known as the heteroscedastic two-limit tobit model in the econometrics literature — are special cases of extended-support beta regression, depending on whether a single extra parameter is zero or infinity, respectively. To overcome identifiability issues due to the similarity of the beta and normal distributions for certain parameter values, we shrink towards beta regression by letting the extra parameter have an exponential distribution with unknown mean. A Gauss-Laguerre quadrature approximation results in efficient likelihood-based estimation and inference procedures, which the betareg R package implements since version 3.2.0. We use XBX regression to analyze investment decisions in a behavioural economics experiment about the occurrence and extent of loss aversion. In contrast to standard approaches, we capture both the probability of rational behaviour and the mean of loss aversion. Extensive comparisons with alternative approaches illustrate the effectiveness of the new model.
@@ 1.1 Bounded responses and beta regression
From the available approaches, beta regression (Ferrari and Cribari-Neto, 2004; Smithson and Verkuilen, 2006) has emerged as the dominant likelihood-based approach for modelling bounded responses in terms of covariates. It has been successfully applied in diverse fields, including biology, ecology, economics and social sciences, finance, manufacturing, genetics, and engineering; see, for example, Allenbrand and Sherwood (2023), Leng et al. (2021), Namin et al. (2020), and Wang et al. (2015) for a few noteworthy applications, and Douma and Weedon (2019) and Geissinger et al. (2022) for recent articles making a case for its use in natural sciences and associated applied fields. The reason for its popularity is that beta regression respects the bounded range of the responses, and allows their mean and dispersion to be expressed in terms of covariates through separate regression structures. Such a construct offers high flexibility in the shape of the response distribution, capturing both the heteroscedasticity and the asymmetries observed in bounded responses.
@@
where $\mu\in(0,1)$, $\phi>0$, $B(p,q)$ $(q,p>0)$ is the beta function (see, Abramowitz and Stegun, 1964, §6.2.1), and $I(A)$ is the indicator function taking value $1$ if $A$ is true, and zero, otherwise. In this parameterization, $\mathop{\rm E}(Y\mid\mu,\phi)=\mu$ and $\mathop{\rm var}(Y\mid\mu,\phi)=\mu(1-\mu)/(1+\phi)$ so that $\phi$ is a precision parameter. The beta regression model assumes that the observed responses $y_{1},\ldots,y_{n}$ are realizations of beta random variables $Y_{1},\ldots,Y_{n}$ that are independent conditionally on explanatory vectors $x_{1},\ldots,x_{n}$ and $z_{1},\ldots,z_{n}$, where $x_{i}$ and $z_{i}$ are $p$- and $q$-dimensional vectors, respectively, possibly sharing some or all components. As is apparent, we are assuming here that the observed responses take value between zero and one or have been transformed to do so as $y_{i}=(y_{i}^{\prime}-a)/(b-a)$, where $y_{i}^{\prime}$ are the original responses and $a$ and $b$ are the minimum and maximum value they can take. The mean and precision of $Y_{i}$ are then linked to the explanatory variables according to
@@
The parameter vectors $\beta$ and $\gamma$ are typically estimated by maximum likelihood, and standard inferential procedures, such as likelihood ratio, Wald, and score tests, apply on the grounds of central limit theorems for the derivatives of the log-likelihood under assumptions on model adequacy. Grün et al. (2012) provide more details and describe a range of modelling strategies and methods based on the beta regression model.
@@ 1.2 Boundary observations
One important limitation of the beta regression model as defined by (1) and (2) is that likelihood-based estimation and inference is not possible when at least one of the observed responses is either zero or one. Specifically, depending on the parameter values, the density (1) at those observations is either zero or infinity. As a result and regardless of the sample size, the log-likelihood of the beta regression model is either zero or infinite or undetermined, and maximum likelihood estimation and the associated inferential procedures cannot be used. This is the reason that popular software for beta regression modelling, such as the betareg R package (Cribari-Neto and Zeileis, 2010; Grün et al., 2012) prior to version 3.2-0, returned an error when there is at least one response value that is zero or one.
One popular and straightforward approach to deal with boundary observations is by Smithson and Verkuilen (2006), who propose to transform the response to $(y_{i}+u)/(1+2u)$, with $u=1/\{2(n-1)\}$, and then fit the beta regression model in (1) and (2) with standard software. The transformed responses, then lie in $(0.5/n,1-0.5/n)\subset(0,1)$, and, as a result, there are no issues when fitting the beta regression model. The proposal in Smithson and Verkuilen (2006) has the advantage of preserving the parsimony and the convenience in estimation and inference that is provided by beta regression model in (1) and (2), but the value of $u$ is ad hoc and, as Example 1 demonstrates, can have a marked effect on inference.
@@
Smithson and Verkuilen (2006, Example 3) models the contribution of nonverbal IQ and dyslexic versus nondyslexic status to the scores of 44 children on a test of reading accuracy. The scores were rescaled to lie in the unit interval, and 13 out of 44 children achieved a score of one. We transform the responses using $u=1/86$, and fit the beta regression model with $g_{1}(\mu_{i})=\log\{\mu_{i}/(1-\mu_{i})\}$, $g_{2}(\phi_{i})=\log(\phi_{i})$ in (2), where $\eta_{i}$ involves an intercept and the main and interaction effects of IQ and dyslexia status, and $\zeta_{i}$ involves only the main effects of IQ and dyslexia status. [^1] The numerical figures in Smithson and Verkuilen (2006, Table 5) can be reproduced by transforming only the boundary responses, and leaving the non-boundary ones unaltered. The same is true with Smithson and Merkle (2013, Table 6.2). We focus on the evidence against the omission of the interaction term.
Figure 1: Likelihood ratio and Wald statistics for the inclusion of an interaction term between dyslexia status and IQ in the setting of Example 1 versus $u$ ($\log_{10}$ scale). The vertical lines are the quantiles of the $\chi^{2}_{1}$ (likelihood ratio test) and normal distribution (two-sided Wald test) at levels $1.0\%$, $2.5\%$, $5.0\%$, $10.0\%$. The vertical dashed line is at $u=1/86$ as per the proposal in Smithson and Verkuilen (2006).
@@
$\mathop{\rm E}(Y)=P(Y^{*}\geq 1\mid\theta)+P(0<Y^{*}<1\mid\theta)\mathop{\rm E}(Y^{*}\mid 0<Y^{*}<1,\theta)\,.$ (3)
@@
where $\prod_{A}a$ is $1$ if $A$ is the empty set. The standard approach towards defining a censored regression model is to put a regression structure on the components of $\theta$. The most widely-used censored regression model is the censored normal one, also known as the two-limit tobit. This model assumes that $Y^{*}_{1},\ldots,Y^{*}_{n}$ are independent conditionally on covariate vectors $x_{1},\ldots,x_{n}$, with $Y^{*}_{i}$ distributed according to a normal distribution with mean $x_{i}^{\top}\beta^{*}$ and variance $\sigma_{i}^{2}=\sigma^{2}$ (see, for example Rosett and Nelson, 1975; Maddala, 1983, Section 6.7, for definition and discussion). The heteroscedastic version of the two-limit tobit assumes that the variances are linked to explanatory variables as $g(\sigma_{i})=z_{i}^{\top}\gamma^{*}$, where $g(\cdot)$ is a monotone link function that maps $(0,\infty)$ to the real line (see, for example, Messner et al., 2016). Dixon and Sonka (1982) compare least squares and maximum likelihood estimation of the two-limit tobit in terms of estimates in an empirical application, and Hoff (2007) compares the two-limit tobit model to the semi-parametric approach in Papke and Wooldridge (1996) and the inflated beta distribution in Cook et al. (2008) for data envelopment analysis. Dixon and Sonka (1982), Hoff (2007), and Greene (2011, §19.3) touch on the interpretation of the parameters of the two-limit tobit. The disadvantage of the two-limit tobit is that it is less flexible than the beta distribution and cannot accommodate features that often occur in bounded-domain variables such as high skewness or kurtosis.
@@ 1.3 Our contribution
Finally, we present a case study highlighting the benefits of XBX regression. We revisit the analysis of an experiment from behavioural economics about the occurrence and extent of loss aversion (Glätzle-Rützler et al., 2015), which was originally analyzed using linear regression for the mean loss aversion only. Here, we demonstrate that the XBX model leads to qualitatively the same results for the effects of mean loss aversion, simultaneously allowing for economically relevant insights about the probability of rational behaviour.
The betareg R package, since version 3.2-0, provides methods for fitting and drawing inferences from XBX regression models.
Section 2 presents the XB distribution and shows that the beta and the censored normal distribution are its limiting cases. Section 3 introduces the XBX regression model, and discusses parameter estimation and inference using approximate likelihood. Section 4 illustrates the benefits of XBX regression through the analysis of a loss aversion experiment from behavioral economics. Section 5 presents a comprehensive simulation experiment comparing XBX regression to the heteroscedastic two-limit tobit model based on the continuous ranked probability score. We conclude with remarks and directions for further research in Section 6.
@@ 2 Extended-support beta distribution
    If $\mu_{*}=\mathop{\rm E}(Y\mid\mu,\phi,u)\in\Re$ and $\sigma_{*}^{2}=\mathop{\rm var}(Y\mid\mu,\phi,u)\in(0,\infty)$ are fixed, then, as $u\to\infty$, $Y$ converges in distribution to $\mu_{*}+\sigma_{*}Z$, where $Z$ is a standard normal random variable.
@@
$h(t\mid\mu,\phi,u)=\frac{\sigma(\mu+\sigma t)^{\alpha-1}(1-\mu-\sigma t)^{\beta-1}}{B(\alpha,\beta)}I\left(\frac{-u-\mu_{*}}{\sigma_{*}}<t<\frac{1+u-\mu_{*}}{\sigma_{*}}\right)\,.$ (6)
@@
A direct corollary of Theorem 1 is that the XB distribution coincides with the beta distribution for $u=0$, and that it converges to the normal distribution censored in $[0,1]$ as $u\to\infty$. Hence, regression models based on the XB distribution allow to parameterize the transition from beta regression to the heteroscedastic two-limit tobit, at the expense of a single extra parameter $u$. The beta distribution and the normal distribution limits of the XB distribution can assume similar shapes in $(0,1)$ for certain parameter settings. Hence, it is natural to consider shrinking $u$ to 00, or, equivalently, shrinking the XB distributions towards the beta distribution, to avoid any identifiability issues that may arise in the estimation of the parameters of the XB distribution, while preserving the flexibility offered by the beta distribution.
@@ 3.1 Exponential extended-support beta mixture
where $f_{\rm(XB)}(y\mid\mu,\phi,u)$ is defined in (7). By its definition, for fixed $\phi$ and $\nu$, the XBX density is symmetric in that $f_{\rm(XBX)}(y\mid\mu,\phi,\nu)=f_{\rm(XBX)}(1-y\mid 1-\mu,\phi,\nu)$. Density (8) is not available in closed form, but it can be accurately and efficiently approximated using a Gauss-Laguerre quadrature rule (see, Abramowitz and Stegun, 1964, Equation 25.4.45) as
@@
Figure 3: Densities of the beta and XBX distribution for all combinations of $\mu\in\{0.5,0.7,0.9\}$, $\phi\in\{0.5,2,20\}$, and $\nu\in\{0.01,0.1,1\}$. The horizontal dotted line is at 1. The height of the vertical black segments represents the point mass at 00 and $1$ for the corresponding parameter setting.
@@ 3.2 XBX regression
From expressions (10), the parameters $\mu$ and $\phi$ have direct interpretations in terms of location and precision of the distribution of $Y^{*}$, as they do in beta regression. In particular, the expectation of $Y^{*}$ is a linear function of $\mu$, with slope $(1+2\nu)$ and intercept $-\nu$. On the other hand, the variance of $Y^{*}$ decreases as $\phi$ increases, with the rate of decrease increasing quadratically with $\nu$.
So, an XBX regression with mean and dispersion effects for the latent distribution can directly be defined by linking $\mu$ and $\phi$ to covariates as in (2). From Figure 3, it is directly apparent that the response distribution in XBX regression can have much more flexible shapes than the censored normal distribution underlying the heteroscedastic two-limit tobit; see, also Figure LABEL:fig-supp:xbx-vs-cn and Figure LABEL:fig-supp:xbx-vs-cn-hist in Section LABEL:sec-supp:xbx-vs-cn of the supplementary material document for a side-by-side comparison of the XBX distribution to the doubly-censored normal distribution.
The expectation and variance of a random variable with XBX distribution are, of course, nonlinear functions of $\mu$, $\phi$ and $\nu$ as they depend on the probabilities of $Y^{*}$ being greater or equal to $1$ and $Y^{*}$ being in $(0,1)$, and on moments of the truncated distribution of $Y^{*}$ in $(0,1)$. Figure LABEL:fig-supp:xbx-e and Figure LABEL:fig-supp:xbx-v of Section LABEL:sec:xbx-e-v the supplementary material show the expectation and variance of the XBX distribution as a function of $\mu$ for a range of values for $\phi$ and $\nu$. For reference, Figure LABEL:fig-supp:xbx-e and Figure LABEL:fig-supp:xbx-v also show the expectation ($\mu$) and variance ($\mu(1-\mu)/\phi$) of the beta distribution as functions of $\mu$ for a range of values of $\phi$.
As expected, for small $\nu$, the expectation and variance of the XBX distribution are almost identical to those of the beta distribution. Notably, for all values of $\nu$ and $\phi$, the expectation of the XBX distribution is found to be a monotone increasing function of $\mu$ and rotationally symmetric at $(0.5,0.5)$. Hence, any interpretations of the signs of the coefficients in the regression structure for $\mu$ in XBX regression are as in standard beta regression, with the usual caveat that changes in covariates in the regression structure on $\mu$ impact also the variance of the response distributions. Like for the beta distribution, the variance of the XBX distribution as a function of $\mu$ is bounded from above for any $\phi$ and $\nu$, with the difference that it can assume bimodal shapes. Furthermore, as for beta, the variance of the XBX distribution decreases as $\phi$ increases. In contrast to beta, however, the mean of the XBX distribution increases with $\phi$ if $\mu>0.5$ and decreases if $\mu<0.5$, most notably for large values of $\nu$. So, in XBX regression, if a regression structure on $\phi$ is present alongside a regression structure on $\mu$, the associated covariates would also contribute to the mean of the response distribution. The same holds for other censored regression models, like the two-limit tobit, where regression structures on the variance of the latent variable, contribute to the mean of the response distribution.